OpenStudy (anonymous):

HELP PLEASE FAST MEDAL/FAN THANKS! Solve the equation by completing the square. x² + 6x = -8

OpenStudy (imstuck):

Take half the x term of 6, which is 3, square the 3 which is 9 and add 9 to both sides of the equation. LIke this:\[x ^{2}+6x+9=-8+9\]This gives you\[(x+3)^{2}=1\]

OpenStudy (imstuck):